OpenStudy (anonymous):

Which polynomial is a difference of two squares? x2 − 16 x2 + 16 x2 − 8 x2 + 8 The third one? @amistre64

OpenStudy (amistre64):

well, you got difference right, its subtraction; but 8 is not a perfect square.

OpenStudy (jhannybean):

\[\large a^2-b^2 = (a-b)(a+b)\]

OpenStudy (anonymous):

Then,the first one. :

OpenStudy (luigi0210):

Here are perfect squares: 1*1=1 2*2=4 3*3=9 4*4=16 5*5=25 6*6=36 7*7=49 8*8=64 9*9=81 10*10=100 I don't wanna go on..

OpenStudy (amistre64):

first one is good :) if we run thru the perfect squares: 0^2 = 0 1^2 = 1 2^2 = 4 3^2 = 9 4^2 = 16 <-- that works fine 5^2 = 25 6^2 = 36